OpenStudy (anonymous):

One more for good measure? @jim_thompson5910 Find the exact value of the radical expression in simplest form. the square root of s to the eighth power + the square root of 25 times s to the eighth power + 2 the square root of s to the eighth power − the square root of s to the fourth power

OpenStudy (anonymous):

\[\sqrt{s^8} + \sqrt{25s^8}+2\sqrt{s^8}-\sqrt{s^4}\]

OpenStudy (anonymous):

The things that look like 8's are s's

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

the first thing to do is to split up \[\large \sqrt{25s^8}\] into \[\large \sqrt{25}*\sqrt{s^8}\]

jimthompson5910 (jim_thompson5910):

what is the square root of 25?

OpenStudy (anonymous):

Five

jimthompson5910 (jim_thompson5910):

so \[\large \sqrt{s^8} + \sqrt{25s^8}+2\sqrt{s^8}-\sqrt{s^4}\] becomes \[\large \sqrt{s^8} + 5\sqrt{s^8}+2\sqrt{s^8}-\sqrt{s^4}\]

jimthompson5910 (jim_thompson5910):

now how do we simplify things like \(\large \sqrt{s^8} \) and \(\large \sqrt{s^4} \) ??

OpenStudy (anonymous):

I forgot

jimthompson5910 (jim_thompson5910):

what is the square root of x^2

OpenStudy (anonymous):

x? I'm not really good with simplifying radical expressions, I had a mod get frustrated with me yesterday because I was terrible at it.

jimthompson5910 (jim_thompson5910):

its ok, take it at your pace, not the mods

jimthompson5910 (jim_thompson5910):

square roots undo squaring, so yes, \(\large \sqrt{x^2} = x\)

jimthompson5910 (jim_thompson5910):

this is assuming that x is not negative

jimthompson5910 (jim_thompson5910):

in general, the rule is this \[\large \sqrt{x^n} = x^{n/2}\]

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

what that means is that \[\large \sqrt{s^8} = s^{8/2} = s^4\] and \[\large \sqrt{s^4} = s^{4/2} = s^2\]

OpenStudy (anonymous):

Would the answer be \[28s^4\sqrt{s}-s^2 ?\]

jimthompson5910 (jim_thompson5910):

notice how \[\large \sqrt{s^2} = s^{2/2} = s^1 = s\] so the rule holds up to the fact that square rooting a square undoes it

jimthompson5910 (jim_thompson5910):

so using those two facts, we can say \[\large \sqrt{s^8} + 5\sqrt{s^8}+2\sqrt{s^8}-\sqrt{s^4}\] turns into \[\large s^4 + 5s^4 + 2s^4 - s^2\] then you combine like terms to get \[\large 8s^4 - s^2\]

jimthompson5910 (jim_thompson5910):

so you were close in a way

OpenStudy (anonymous):

Oopsies, alright. :o ^_^ Thanks again.

jimthompson5910 (jim_thompson5910):

sure thing